OpenStudy (anonymous):

Evaluate: -6 - (-3) PLease HELPPPPP !!!

OpenStudy (anonymous):

the answer is -3

OpenStudy (solomonzelman):

\(\large\color{blue}{ \sf -6-(-3)=-6+3=-3}\)

OpenStudy (solomonzelman):

minus minus = plus

OpenStudy (anonymous):

OMG thanks so much guys <3 #lifesavers

OpenStudy (anonymous):

Sandy cut out 2 squares and 7 circles from a piece of cardboard. The length of the sides of each square was x inches and radius of each circle was y inches. The expression below shows how many more square inches of cardboard Sandy used for making the circles than the squares. 7πy2 - 2x2 What does the term 7πy2 in the expression represent?

OpenStudy (solomonzelman):

The expression is 7πy^2 - 2x2. ( Not 7πy2 - 2x2 ) y is the radius, so πy^2 (is just the same as πr^2 ) saying that πy^2 is the area of each circle or the amount of area of inches cut to make circles. So it represents the amount of cardboard used on circles.

OpenStudy (solomonzelman):

And the seven in "7πy^2" is the number of circles.

OpenStudy (anonymous):

omg thanks<3

OpenStudy (solomonzelman):

You welcome !